Given teh function H described by h(x)=6x, find each of the following:
a) h(-3)
b) h(2)
c) h(28)

Answer by checkley77(12844)

h(x)=6x
h(-3)=-3*6=-18
h(2)=2*6=12
h(28)=28*6=168